40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog, Page 2 of 56  Not logged in ELOG logo
ID Date Author Type Category Subject
  2741   Tue Mar 29 09:13:20 2022 AidanSafetyCleanlinessLab flooding

Facilities placed a blower and dehumidifier in B265B. I checked the airflow and the air around the tables is comparitively still. The North table is covered and the South table is over pressurized by HEPA filters, so there should be little risk of dust being stirred up.

Quote:

Flood photo album: https://photos.app.goo.gl/BZAG8DyQzFVTfMNz6 (This link is read-only who has no access to the account)

 

 

Attachment 1: IMG_8104.jpg
IMG_8104.jpg
Attachment 2: IMG_8103.jpg
IMG_8103.jpg
Attachment 3: IMG_8102.jpg
IMG_8102.jpg
Attachment 4: IMG_8101.jpg
IMG_8101.jpg
Attachment 5: IMG_8100.jpg
IMG_8100.jpg
  2740   Mon Mar 28 18:00:43 2022 KojiSafetyCleanlinessLab flooding

Flood photo album: https://photos.app.goo.gl/BZAG8DyQzFVTfMNz6 (This link is read-only who has no access to the account)

 

  2739   Mon Mar 28 16:09:57 2022 AidanSafetyCleanlinessLab flooding

Some photos of affected areas in B265A and B265B (elog shows some preview photos - click on PDF for full set).

Stephen did a great job cleaning up and drying up. Most equipment is powered off and we're leaving it off for a couple of days to dry completely. We'll want to check the stuff on the red lab cart thoroughly. 

Quote:

When I went into QIL today there was a lot of flooding from water dripping from the ceiling at several places in the lab. Images attached.

 

Attachment 1: Flood_in_sub-basement-20220328.pdf
Flood_in_sub-basement-20220328.pdf Flood_in_sub-basement-20220328.pdf Flood_in_sub-basement-20220328.pdf Flood_in_sub-basement-20220328.pdf Flood_in_sub-basement-20220328.pdf Flood_in_sub-basement-20220328.pdf Flood_in_sub-basement-20220328.pdf Flood_in_sub-basement-20220328.pdf
  2738   Mon Mar 28 14:29:22 2022 shrutiSafetyCleanlinessLab flooding

When I went into QIL today there was a lot of flooding from water dripping from the ceiling at several places in the lab. Images attached.

Attachment 1: Flooding.zip
  2737   Sat Mar 26 12:49:18 2022 StephenDailyProgressCryo vacuum chamberWarmup started, issue with steady state heating

25 March 2022 (Friday) at 21:00, went to QIL to start warmup.

 - Cryocooler was turned off at 21:21

 - Heater output was disabled - it seemed there was an issue, and therefore I opted for passive warmup only.

 

Heater Issue Troubleshooting

Symptom: Heater Output was enabled but reporting only .35 W and "Err" indicator.

Symptom: When output was disabled, a fan noise was terminated. When output was reenabled, the same fan kicked back on. The fan was driving much harder than I had ever heard it before.

Symptom: The output indicated .35 W but the test mass temperature was 66 K. Past heater power for steady state at 120 K was on the order of 1 W.

Per CTC 100 manual:

   - pg 9 (100W heater outputs) indicates: "If the temperature of either PCB exceeds 60°C, the CTC100 automatically shuts off the corresponding output" which was not the case.

   --> Apparently not an overtemperature situation.

   - pg 9 (Hardware faults) indicates a list of error conditions which are accompanied by pop up windows.

   --> This error had no pop up window, not quite sure what to make of that except that the controller doesn't think our issue is something it can identify.

   - pg 29 (The system fan) notes that "The main system processor reads the desired fan speed from each I/O card and sets the fan to the fastest requested speed".

   --> Suggests that the louder fan noise may have indicated higher temperature condition, even if not an over tempearature condition.

   - pg 41 (Numeric) describes that in the typical numerical view of the data channels, the message "Err" that I saw on the heater channel indicates "an internal error has occurred".

   --> No explanation of what an "internal error" is, but in this case I suspect it could reflect that the heater output is not coupled the input Workpiece temperature.

Best Guess: the symptoms and the lack of any apparent controller-identified fault suggests that the heater may have debonded. I didn't look at temperature history, so I'm not sure if there was a point where the heater was bonded to the test mass during this run.

Next Steps: We should open up and investigate.

  2736   Thu Mar 24 16:08:35 2022 RadhikaDailyProgressCryo vacuum chamberMegastat cooldown 3/18 - adding indium gaskets to copper bar joints

The data from this cooldown is attached (labeled 03/19 - UTC time), compared to the run started on 03/10. In between these 2 cooldowns, the greased joints were replaced with indium joints on both sides of the copper bars (cold head to copper bar, copper bar to flexible strap). 

Observations:
1. The outer shield is in weaker conductive contact with the cold plate than before. Due to the resting of the outer shield (no clamping), the amount of contact is expected to vary run to run.
2. The cold head seems to be experiencing a larger heat load than compared to the previous run. The inner shield does get slightly colder, but this could be attributed to less cooling power being diverted to the outer shield. It is unclear what additional elements could be receiving this extra cooling power.
3. The steady-state temperature difference between the cold head and inner shield is much smaller than in previous runs (23K to 10K). Since this offset is determined by steady-state heat loads along the conductive pathway to the inner shield, it isn't immediately clear why this would have decreased
4. The test mass cooldown trend stays nearly identical to the previous run, despite the aforementioned differences. This seems to confirm that there is some heat leaking from the outer shield to the test mass, which balances the extra cooling power from the inner shield. 

Efforts to update the model (indium links) and analyze these runs is ongoing. Accurate analysis rests on understanding points 2 and 3. above, since the current model predicts a much larger steady-state offset between the cold head and inner shield.

I plan to devote some time to this analysis before planning another Megastat cooldown.

Attachment 1: cooldown_0319_vs_0310.pdf
cooldown_0319_vs_0310.pdf
  2735   Tue Mar 22 09:19:38 2022 shrutiUpdateWOPOWaking up WOPO - back to green path

[Yehonathan, Shruti]

Yesterday we went back to fiddling with the green path. Soon after opening the green shutter and then switching the doubling cavity to 'AUTO' we were able to see 150 mW of green light. We were able to replicate this a couple of times yesterday.

Since we had earlier removed the green fiber from the fiber launch to clean its tip, the coupling into the fiber turned out to be quite poor. As can be seen in Attachment 1, Yehonathan pointed out that a lot of green light was being lost to the cladding due to poor coupling. He then played around with the alignment and finally was able to see 65% coupling efficiency. This process seemed to involve a great amount of trial and error through several local power minima.

Attachment 2 shows that the coupling between the two fibers at the 532 nm input of the waveguide is quite poor (there is visible light being lost in the cladding). Furthermore, this light intensity decreases as we get closer to the waveguide meaning this light is being dissipated in the fiber. Even at the 1064 nm output where we expect to see squeezing there is some remnant green light.

We wanted to test if the green leakage reaching the PDs were causing additional noise. For this we just looked at the spectrum analyzer on the Moku (after amplifying 100x with the SR 560) and saw no difference in the noise spectrum with and without the green shutter being open. Although, we're not convinced with this measurement since we were not able to find good quality SMA cables for the entire path. Moving around the BNCs seemed to change the noise. Also, near the end, we noticed some coupling between the two channels on the Moku while measuring the noise that seemed to cause additional noise in one of the channels. We did not have sufficient time yesterday to probe this further.

Attachment 1: before_opt.jpg
before_opt.jpg
Attachment 2: around_waveguide.png
around_waveguide.png
  2734   Fri Mar 18 19:34:35 2022 RadhikaDailyProgressCryo vacuum chamberMegastat cooldown 3/18 - adding indium gaskets to copper bar joints

Today I opened up Megstat to add indium in between the copper bar joints, with the hopes of speeding up cooldown and informing the thermal model.

Outline of procedure:

  • Disconnected RTD pins/sockets and removed suspension from chamber (with test mass)
  • Removed inner shield and outer shield from chamber       
    • Cleaned up residual grease/aquadag from cold plate
  • Opened up bottom conflat of the T to access ends of the 2 copper bars
  • Unbolted copper bars and removed from chamber [Attachment 1]
    • Gave copper bars a good wipe down with isopropyl to remove old grease
  • Placed 2 indium gaskets previously prepared by Stephen between joints of the two bars [Attachments 2, 3, 4]
    • Fed the bars back through the chamber tube and placed additional indium gasket between top bar and cold head
    • Placed last indium gasket between flexible strap and bottom copper bar in chamber body
  • Bolted up the copper bars, passed through 2 sheets of aluminized mylar to insulate copper bars [Attachment 5]
  • Resecured the cold head RTD with spring clamp, inserted mylar "cap" around cold heat [Attachments 6, 7]
  • Touched up shields
    • Reapplied grease to bottom lip of inner shield
    • Added some more kapton tape to secure peek sheets to bottom lip of outer shield
    • Reinserted shields, bolted down inner shield
  • Reinserted suspension frame with test mass [Attachment 8]
  • Reconnected RTDs and closed up

The roughing pump was turned on at 7:20pm, followed by the cryocooler at 7:50pm.

Attachment 1: IMG_3191.jpeg
IMG_3191.jpeg
Attachment 2: IMG_3197.jpeg
IMG_3197.jpeg
Attachment 3: IMG_3194.jpeg
IMG_3194.jpeg
Attachment 4: IMG_3195.jpeg
IMG_3195.jpeg
Attachment 5: IMG_3206.jpeg
IMG_3206.jpeg
Attachment 6: IMG_3204.jpeg
IMG_3204.jpeg
Attachment 7: IMG_3205.jpeg
IMG_3205.jpeg
Attachment 8: IMG_3210.jpeg
IMG_3210.jpeg
  2733   Wed Mar 16 12:22:44 2022 YehonathanUpdateWOPOWaking up WOPO - attempts at readout

{Shruti, Yehonathan}

Yesterday, we measured a bunch of noises.

We wanted to have as reference the Moku noise, the PDs noise, and measure the shot noise of the LO again.

Attachment 1 shows the Moku noise measured by just taking data with no signal coming in. We tried both the spectrum analyzer (SA) and the oscilloscope tools, with and without averaging, and the difference between the channels.

For some reason, the SA has a worse noise figure than the oscilloscope and the difference channel doesn't give us any special common-mode rejection. Also more averaging doesn't help much because we are already taking 1.2ms of data which is way longer than 1/RBW=0.2ms we are taking here.

From now on we use the oscilloscope as the spectrum analyzer and to its noise we refer as the Moku noise floor.

Moving on, we try to measure the PD dark noise. Given that the PD dark noise floor is ~ 6nV we don't expect to see it with the Moku without amplification. Attachment 2 shows that indeed we couldn't resolve the PD dark noise.

We then opened the LO shutter. We measured with a power meter 1mW and 1.15mW coming impinging on the PDs. The voltage readings after the preamp were 1.66V for the white fiber, and 1.93 V for the red fiber. These values suggest responsivities of 0.830 and 0.834 respectively.

The PDs were measured using the Moku scope and subtracted digitally with some small gain adjustment (0.93*ch1-1.07*ch2) between the channels. The result is shown in attachment 3 together with the expected shot noise level.

1. There is not enough clearance for detecting squeezing.

2. Expected shot noise level is still too high. Does the 2kohm preamp gain go all the way above 1MHz??

Attachment 1: Moku_Noise.pdf
Moku_Noise.pdf
Attachment 2: PD_Connected_no_light.pdf
PD_Connected_no_light.pdf
Attachment 3: Diff_Channel.pdf
Diff_Channel.pdf
  2732   Mon Mar 14 14:04:54 2022 RadhikaDailyProgressCryo vacuum chamberInsulated outer shield cooldown started 3/10

The heater was turned on at 2:05PM 3/14, with a setpoint of 123K.
The cryocooler was turned off at 10:50AM 3/15, and the heater setpoint was raised to 275K to aid in warmup.

  2731   Thu Mar 10 17:12:01 2022 awadeUpdateWOPOWaking up WOPO - attempts at readout

[awade]

Good to see this experiment being revived.

1. The design of this laser had a number of flaws and one of them is this sensitivity to backreflections at 532 nm. I mostly just disabled the doubler's lock and closed the shutter for good measure, but probably best not to leave flickering around in an unstable state when you're away.

2. I built in the inversion in the second channel to give myself the option to electronically subtract: something that didn't end up being very practical compared to just digitally recording channels and subtracting in post.

  • I'm surprized the SR560 don't given you clearance there.  Nominally these units are 4 nV/rtHz if you operate in low noise mode AND gain=100 (see p21 of the manual), below this gain gives you 10-60 nV/rtHz noise. When I built the PD circuits I did verify that I was getting the clearanceI expected.  At 1 mW on each photodiode one would expect order =sqrt(2*h*c/lambda0*Responisvity*Power)*Gain ~ sqrt(2*6.626e-34*3e8/1064e-9*1e-3*0.7)*2e3 = 32 nV/rtHz.  
  • A few things to verify:
    • check the DC voltage (just with a multimeter) to see true power picked up by diodes, this should be 1.4 V for about 1 mW of 1064 nm;
    • Make sure you're AC coupled into SR560, there is no way you operate at gain 100 or above and also not saturate for 1 mW (~1.4 V amp output DC) of light
    • at 2kΩ gain you should expect the noise floor to be of order =sqrt(4*kB*T*G)=sqrt(4*1.38e-23*300*2e3)~5.8 nV.  Only just clear of the SR560 spec, and about equal to typical actual performance.  To see this level you might want to pre-amplify with a Femto amplifier, the 40m Busby box or the ganged amplifier box I made for the CTN lab (its black with gold writing, Anchal knows the one). A dark measrument like this may have a little offset that you can either null or just AC couple with a minicircuits DC block;
    • Take a terminated (50 Ω) measurment of your ADCs when you collect your PD 'dark' data. Even better also collect terminated SR560 data. And put these on the plot.  Moku:Labs have ~30 nV/rtHz @100 kHz and above.  Just be sure you're measure photodetectors and not Pre-amp or ADC noise. Moku nominal input refered noise is 13.9 nV/rtHz * sqrt(1+220kHz/f).
    • If you can't get any quick progress, try all the above with minicircuits lower noise amplifiers.  They have plenty of bandwidth and go to higher frequencies.
  • Just measureing the output of the PD directly, with no subtraction or amplification, I'd say you are looking at laser technical noise at about 1 MHz: this is what the subtraction is for, to null the LO noise effects to only listen to signal port.  Maybe somethings burried in the subtracted signal offline, but you need some simultaneous termianated measurements back long the signal chain to put some bonds on what is the limiting noise here.

3. Subtracted noise spectra

  • These gain numbers sound right to me.
  • The AD829 is designed to drive this combined 150 Ω load, I would stick to 50 Ω terminations for now. On the topic of mokus: again verify its input referred noise and pre-amplify accordingly.  Also there is a choice of "Normal" or "Precision" acquisition mode, I think the right choice is precision (this should have the right filters to kill aliasing from the downsampling)
  • ~70 nV/rtHz shot noise sound about right to me.  Not clear why a subtracted signal doesn't seem to reach this.  Once again, measure the actual DC voltage output from the TIA to get the true absorbed photons and use that to calibrate your estimate.  

We should chat some time on zoom about more details (rana can forward my details).  Hope this enought to go on for at least the homodyne part of the experiment.  

  2730   Thu Mar 10 16:21:11 2022 RadhikaDailyProgressCryo vacuum chamberInsulated outer shield cooldown started 3/10

[Stephen, Radhika]

Our goal for this week's cooldown was to tape peek sheets fully around the outer shield lip, to leave no bare aluminum contact area with the cold plate. Secondly, we wanted to diagnose and mend the issue preventing the heater from outputting any power. The full procedure was:

1. We allowed the chamber to vent, unbolted the chamber lid and outer/inner shield lids.

2. We noticed that the solder joints between the heater body and its leads had debonded [Attachment 1].

    a. The suspension frame was taken out of the chamber and the test mass was removed from the frame.

    b. In doing so, we noticed that the varnish joining 1. the heater to cigarette paper and 2. cigarette paper to Si was debonding in certain areas, likely due to Aquadag not being fully removed from the test mass in the area of contact [Attachment 2].

3. We wrapped the copper leads a few times around the heater "wings" and re-applied solder [Attachments 3, 4].

4. We cleaned off aquadag from a greater area on the test mass and applied varnish to re-bond the heater [Attachments 5, 6].

    a. We let the varnish cure for ~2 days with a small weight on top.

5. The outer shield was removed from the chamber (without unbolting/removing the inner shield), and a single layer of peek sheet was taped the whole way around the bottom lip [Attachment 7].

6. We re-inserted the outer shield and passed the RTDs back through.

    a. We reattached a few RTD lead pins/sockets that had broken off in handling.

7. Lastly, we placed the test mass back into the suspension and into the chamber.

8. Close out [Attachments 8, 9]

The vacuum pump was engaged and the cryocooler was turned on at ~3:30PM.

Attachment 1: 82E812D8-946F-4F93-95C2-A62EC189A1E5_1_105_c.jpeg
82E812D8-946F-4F93-95C2-A62EC189A1E5_1_105_c.jpeg
Attachment 2: CFDD543D-17FB-4F49-A62B-090A15E47821_1_105_c.jpeg
CFDD543D-17FB-4F49-A62B-090A15E47821_1_105_c.jpeg
Attachment 3: B5D683AC-986A-49B9-A1C3-379851C877E5_1_105_c.jpeg
B5D683AC-986A-49B9-A1C3-379851C877E5_1_105_c.jpeg
Attachment 4: E06C8191-8A7F-4917-A97A-80F9F38BA614.jpeg
E06C8191-8A7F-4917-A97A-80F9F38BA614.jpeg
Attachment 5: 59F4BAD9-C23E-4881-AC82-AEADD05B3D45.jpeg
59F4BAD9-C23E-4881-AC82-AEADD05B3D45.jpeg
Attachment 6: IMG_3160.jpeg
IMG_3160.jpeg
Attachment 7: C8A0E8EB-EB18-4B76-BF0D-6A67A4831C77_1_105_c.jpeg
C8A0E8EB-EB18-4B76-BF0D-6A67A4831C77_1_105_c.jpeg
Attachment 8: IMG_3167.jpeg
IMG_3167.jpeg
Attachment 9: IMG_3168.jpeg
IMG_3168.jpeg
  2729   Fri Mar 4 13:18:09 2022 RadhikaDailyProgressCryo vacuum chamberDouble copper bar cooldown rerun started 02/23

Analysis of 02/24 cooldown data

Attachment 1 shows the cooldown data for this run. Attachment 2 compares this run to the previous 02/11 run, where in between insulating peek sheets were taped to 2 locations along the bottom rim of the outer shield. 

Observations:

  1. The inner shield, outer shield, and test mass all cool slightly faster initially in this run (02/24) compared to 02/11. This effect is seen until ~35 hrs, after which:

  2. The outer shield starts to warm up and re-equilibriate. It seems the radiative heating from the chamber strongly kicked in once the outer shield was sufficiently cold.

The best fit for the data can be seen in Attachment 3. Note the addition of the copper bar model, which considers radiative heating from the chamber at RT. 

Comments:

  1. The outer shield is still getting quite cold, so we have to consider increasing the insulation from peek sheets (either adding more layers or additional points of contact), or another approach altogether.

  2. There are still obscure effects at play in early cooldown that the model is not considering. I have gone back to the drawing board and am trying to fit the raw inner shield data to a sum of exponential terms, in hopes of narrowing down the cooling mechanisms that could be affecting data. 

Next steps:

- Check on the heater leads during next opening and perform tests to ensure test mass is warming up

- Devise insulation solutions for outer shield to decrease system heat load

- Consider using indium foil to increase thermal conductance between joints along cooling pathway

 

Attachment 1: 02-24_cooldown_data.pdf
02-24_cooldown_data.pdf
Attachment 2: 02-24_cooldown_data_comparison.pdf
02-24_cooldown_data_comparison.pdf
Attachment 3: cooldown_0224_bestfit.pdf
cooldown_0224_bestfit.pdf
  2728   Fri Mar 4 11:49:45 2022 shrutiUpdateWOPOWaking up WOPO - attempts at readout

[Yehonathan, Shruti]

1. Doubling cavity and green beam

Since we had left the lasers ON with the shutters closed we wanted to see if the powers measured after opening the shutter would be similar to what it was when we left. We realized that opening and closing the green shutter destabilizes the doubling cavity (the FI is after the shutter and the shutter does not seem to be a good dump), which in turn changes the SHG crystal temperature (possibly because of the power fluctuation within the crystal). Re-opening the shutter requires some tuning of the temperature and offset to recover similar output power. Finally, after some tuning, we were able to see 156 mW of green light.

 

2. Attempt at measuring LO shot noise

  • We want to measure the BHD output A-B channel, which we expect to be dominated by shot noise since all the classical noise would be canceled when optimally balanced, but found that one of the PDs was inverted so that the sum of the two channels would be what we needed to measure. Since the SR 560 has only an A-B option, we used a second SR 560 to invert the B channel before subtracting
  • Operating at an LO power of ~4 mW did not give us sufficient clearance from the dark noise in each channel with the SR 560s, which was around the max power I believe we're supposed to use for the BHD LO
  • Somehow measuring the output directly, without the SR 560, gave us some clearance over the dark noise at ~1 MHz and higher (possibly because 1 MHz is the SR560's BW) so we decided to measure the time series of both channels together and do the optimized subtraction and FFT offline

3. Subtracted noise signal spectra [Attachment 1]

  • The plots show the noise spectra of the channels measured individually. The 'gain adjusted' means that A was multiplied by 1.05 and B by 0.95 in order to get the two plots to more or less line up
  • We used the Moku to measure the timeseries at a sampling rate of 10.2 MS/s for a period of 1.2 ms with AC coupling and 50 Ohm impedance. Elog 2324 suggests the designed measurement was for a 50 Ohm load so we should be impedance matched but I'm yet to convince myself
  • Our estimate for the shot noise was 77 nV/rtHz for 4 mW of power and 87% QE, using a TI gain of 2kOhm (the black dashed line in the plot). If we were impedance matched the yellow trace must be higher than this estimate
  • In our next measurements, we will also record the dark noise, carefully measure the power. There is obviously sonething wrong with the plot

 

 

30 Mar 22 edit: script here, data here

Attachment 1: LO_shot_noise.pdf
LO_shot_noise.pdf
  2727   Wed Mar 2 14:38:55 2022 YehonathanUpdateWOPOWaking up WOPO - some more fiddling and a plan

{Shruti, Yehonathan}

We made some a list of some random questions and plans for the future. We then went down and found answers to some of those:

1. Why is there no Faraday isolator in the 1064nm beam path? (edit: turns out there is, but inside the laser, see pictures in this elog).

2. Do the fibers joined by butt-coupling have similar mode field diameter? If not it can explain many loss issues.

a. In the green path we find that according to the SPDC datasheet the 532nm fiber (coastalcon PM480) is 4um while the input thorlabs fiber (P3-488PM-FC2) coupled to it has an MFD of 3.3um. This mismatch gives maximum coupling efficiency of 96%. Ok not a big issue.

b. At the 1064nm output the SPDC fiber is PM980 with MFD of 6.6um while the BS fiber is 6.2um which is good.

3. What is the green fiber laser damage threshold? According to Thorlabs it is theoretically 1MW/cm^2 practically 250kW/cm^2 for glass air interface. With 3.3um MFD the theoretical damage threshold is ~ 80mW and practically  ~ 20mW. It doesn't sounds like a lot. More so given that we could only get 50% coupling efficiency. How much is needed for observable squeezing? There is the possibility to splice the fiber to an end cap to increase power handling capabilities if needed.

4. Is stimulated Brillouin back scattering relevant in our experiment? According to this rp photonics article not really.

5. How much green light is left after the dichroic mirrors? Is it below the shot noise level? Should check later.

In addition, we found that the green fiber input and the 1064nm fiber output from the SPDC were very dirty! We cleaned them with a Thorlabs universal fiber connector cleaner.

 

 

 

 

  2726   Tue Mar 1 15:59:34 2022 RadhikaDailyProgressCryo vacuum chamberDouble copper bar cooldown rerun started 02/23

The heater was turned on on Tue, 3/1 at 4pm, with control setpoint 123K.

*UPDATE: After checking a few hours later, I noticed the test mass temperature hadn't risen, and the heater power was reading nan. When I initially turned the heater on, I watched the power ramp up to 22W (max power limit) and the test mass temperature start to rise. I wonder if somehow the lead pins shorted after it was turned on. For now I have turned the heater output off and will check on this after warmup.

The cryocooler was turned off at 5:45pm.

  2725   Fri Feb 25 17:09:53 2022 shrutiUpdateWOPOWaking up WOPO - green beam

[shruti, yehonathan]

SHG and 532 nm beam alignment

Yehonathan brought over 532nm/1064nm laser goggles from the 40m.

  • We turned on the 1064 nm Mephisto, and set the pump current initially to 1.5 A (which gave us ~30 mW of output)
  • We then turned on the doubling crystal. It took a while to reach its setpoint temperature of 110 C. Initially (without adjusting the SHG cavity parameters) we saw less than a microW of power after removing the beam block and opening the shutter.
  • Playing around with the SHG cavity settings, we realized that
    • Setting the switch to "Auto" is how to nominally operate the 532 nm beam
    • "Scan" is useful for scanning the cavity, the amplitude of which can be controlled by the "Scan amplitude" knob. "Standby" seems to effectively turn off second harmonic generation
    • "Offset" knob tends to change the amount of green power generated and adjusting the "Gain" simultaneously helps stabilize this power (with some difficulty)
  • On increasing the laser driver current to 2 A and adjusting the temperature setpoint to 109.7 C, we were able to see 100 mW of green power!
  • Next, we played with the alignment into the fiber, seeing that initially there was barely any power at the other end of that patch cable
    • We adjusted the waveplate near the laser head to give us 5.3 mW of power right before coupling into the fiber
    • Adjusting a single mirror (the nearer one) did not result in much gain, so we simultaneously adjusted the two steering mirrors and achieved about a 50% coupling. (Our readings suggested it could be the max)
    • At  the end of the alignment: 2.74 mW at the output and 5.46 at the fiber input
  • Reconnecting the fiber to the waveguide, without adjusting the temperature of the advr waveguide, we saw that the fiber at the output of this crystal seemed to glow.
    • The power measured at the end of the output fiber (fiber after the waveguide) was 0.6 mW. Not entirely sure what the contribution of loss was in the decrease from 2.74 mW through the waveguide.
  • The laser is still ON although the shutters to the green and IR paths are closed. Safety glasses required before opening shutters.

 

Questions about the setup

  1. The spec sheet on the wiki mentioned PM980 and PM480 input and output fibers, respectively for the waveguide operating as an SHG, what were being used instead were P3-1064PM-FC2 and P3-488PM-FC2. Is this a significant source of loss that can be easily remedied?
  2. Yehonathan mentioned that stimulated Brilluoin scattering occurs in all fibers above a threshold. What is the threshold for the the ones used in the setup? We probably want to operate below this threshold.

 

Our next step would be to measure the LO shot noise.

 

 

 

  2724   Thu Feb 24 10:44:06 2022 RadhikaDailyProgressCryo vacuum chamberDouble copper bar cooldown rerun started 02/23

[Radhika, Aaron]

The goals for this cooldown are:

1) Securely re-attach the RTD to the cold head, to fix wacky readings.
2) Reduce thermal conductance between the outer shield and cold plate (/copper bar), to prevent the outer shield from cooling below 200K.
       2a) Stabilize the outer shield so that its contact area with the cold plate is consistent run-to-run. ([Note] Follow-up to an in-person comment from Rana: the positioning of the shields does not leave space for any clamps to secure the outer shield to the cold plate.)

On Tuesday 2/22, we opened up the bottom conflat of the T to check on the RTD spring-clamped to the cold head. I re-inserted the RTD and tightened the nut further than last run, and it seemed much more secure [Attachment 1]. I re-inserted the mylar "cap" covering the cold head [Attachment 2].

In the chamber body, we carefully passed the RTD leads through the inner shield and outer shield apertures to remove the outer shield. We did this without having to unclamp/remove the inner shield or any components inside, to preserve consistency with the last cooldown. A few pins were damaged in this process (from inner shield).

Once the outer shield was removed, we used kapton tape to secure strips of peek sheets to its bottom rim [Attachments 3,4]. The strips were taped at 2 points along the rim associated with the most wobble, with hopes of stabilizing the shield as much as possible.

On 2/23 I repaired the pins previously damaged. I also added kapton tape labels to the socket leads, corresponding to the shapes found on the RTD leads (semi-circle example in Attachment 5). This way it will be much easier to match the right pins and sockets in the future. 

I then bolted up the chamber (close-out pictures can be found on the QIL Google photo dump). The vacuum pump was turned on at 5:45pm, and the cryocooler was turned on at 7:08pm.

Attachment 1: IMG_3095.jpg
IMG_3095.jpg
Attachment 2: IMG_3096.jpg
IMG_3096.jpg
Attachment 3: IMG_3083.jpg
IMG_3083.jpg
Attachment 4: IMG_3087.jpg
IMG_3087.jpg
Attachment 5: IMG_3092.jpg
IMG_3092.jpg
  2723   Tue Feb 22 07:53:06 2022 YehonathanUpdateWOPOWaking up WOPO

{Shruti, Yehonathan}

On Friday, we came down to QIL to poke around the WOPO setup. The first thing we noticed is that the setup on the wiki page is obsolete and in reality, the 532nm light is coming directly from the Diablo module.

There were no laser goggles for 532nm so we turned on the 1064nm (Mephisto) only. The pump diode current was ramped to 1A. We put a power meter in front of Mephisto and opened the shutter. Rotating the HWP we got 39mW. We dialed it back so that 5mW is coming out of the polarizer.

The beam block was removed. We disconnected the LO fiber end from the fiber BS - there is light coming out! we connected a power meter to the fiber end using an FC/PC Fiber Adapter Plate. The power read 0.7mW. By aligning the beam into the LO fiber we got up to 3.3mW.

We connected the BHD PDs to the scope on the table to observe the subtraction signal. Channel 2 was negative so we looked at the sum channel.

Time ran out. We ramped down the diode current and turned off Mephisto.

Next time we should figure out the dark current of the BHD and work toward observing the shot noise of the LO.

  2722   Wed Feb 16 11:35:18 2022 StephenDailyProgressCryo vacuum chamberDoubled Thermal Linkage Capacity run started 11 Feb 2022

[Stephen, Radhika]

The heater was turned on on Wed, 2/16 at 11:30am, with control setpoint 123K. The lower power limit was verified to be 0W.

The cryocooler was turned off on Thu, 2/17 at 12pm. The heater control setpoint was changed to 295K for warmup. The plan is to address the wacky cold head RTD on Monday.

  2721   Tue Feb 15 16:54:27 2022 ranaDailyProgressCryo vacuum chamberDoubled Thermal Linkage Capacity run started 11 Feb 2022

I've been assuming that the inner shield can be treated as a point mass, but perhaps the thinness make a significant delay between the temperature of the cold plate and the inner shield during the initial cooldown.

Could you model the cold shield to estimate what the temperature gradient would look like during the rapid cooldown? Not full 3D, but something approximate that takes into account the conductivity, thinness, and heat capacity.

  2720   Tue Feb 15 11:59:22 2022 RadhikaDailyProgressCryo vacuum chamberDoubled Thermal Linkage Capacity run started 11 Feb 2022

*Note: The RTD spring-clamped to the cold head gave spazzy readings for this cooldown, so the last cooldown's cold head temperature data was used instead for reference.

Looking at the data, there are some initial noteworthy observations:

The outer shield's conductive coupling to the copper bar / cold plate is much higher than previous cooldowns. Attachment 2 shows that the outer shield gets colder than the test mass, and around the 82 hr mark their temperatures cross.
We hoped to see a reduction by about a half in cooldown time for the 02/11 run, after effectively doubling the cross-sectional area of the copper bar. However, the data does not show x2 improvement, as seen in Attachment 1.

It could be that somehow the resistances of re-bolted joints increased significantly to compensate the lowered resistance of the bar, but this doesn't seem too likely. The more likely answer is the model overestimated the original resistance of the bulk of the copper bar relative to other components/joints in the chain. This means more work needs to be done, and hopefully a more realistic model will also resolve the discrepancy in early cooldown of the inner shield data. 

Attachment 2 shows the best fit for the new cooldown.

Attachment 1: Cu_bar_comparison.pdf
Cu_bar_comparison.pdf
Attachment 2: cooldown_0211_all.pdf
cooldown_0211_all.pdf
  2719   Wed Feb 9 15:21:27 2022 ranaHowtoTempCtrlplots

In the plots comparing data and models, can you use the legend to indicate which is which? e.g. use dots for data and solid lines for models, and then label them as that in the legend. Also nice to include error bars on the temperautre measurements; I think there's a python way to plot this as a shaded region as well.

  2718   Mon Feb 7 16:06:36 2022 StephenDailyProgress 31 Jan Fastest Radiative Cooling run, ended 07 Jan

[Radhika, Stephen]

The heater was turned on at 3:13pm on Friday 2/4.

We specified a set temperature of 123K. However, the CTC100 PI control included a 1 W lower limit on the input to the heater, so there was a steady load of 1 W applied to the Silicon Workpiece over the weekend.

At 16:01 the cryocooler was turned off to start the warmup.

The CTC100 PI control was configured with a setpoint of 250 K on the Workpiece RTD, to aid in the warmup, and an allowable power range from 0 W to 22 W.

  2717   Mon Feb 7 11:57:38 2022 StephenDailyProgressCryo vacuum chamberDoubled Thermal Linkage Capacity run started 11 Feb 2022

[Stephen, Radhika]

The double copper bar configuration pictured in Attachment 1 has been implemented. We completed the updates within work sessions on Thursday and Friday. Here's a pseudo-log:

  • Removed all items from coldplate to make space for thermal linkage.
  • Prepared thermal linkage for installation.
    • ref. QIL/2694 for all preparations made.
  • Installed thermal linkage.
    • Removed thermal linkage to allow for regreasing
      • Discovered a single brass #4 washer sandwiched between coldhead and rigid copper bar (likely lost during efforts to install an RTD during intitial thermal linkage install); thermal grease was mostly indented at brass washer and at area opposite location of brass washer, suggesting non-uniform contact.
    • Applied Apiezon N thermal grease at joints: [coldhead - top copper bar], [top copper bar - bottom copper bar, near coldhead], [top copper bar - bottom copper bar, near thermal strap].
    • Sandwiched together top and bottom copper bars, to insert simultaneously.
    • Angled top and bottom copper bars to pass below coldhead, and overall passing the two bars in together went smoothly.
    • Bolted both copper bars into coldhead using long (1.25") 4-40 bolts threaded into coldhead. Access was from bottom port below cryocooler. Torqued using short arm of allen key (no access for long arm).
    • Borrowed one bolt location for RTD spring clamp mount. Confirmed continuity of coldhead RTD.
  • Repaired pins damaged in removal, and inserted sheaths intended to replace the previous scheme and hopefully prevent pin and socket damage. More detail later :).
  • Installed shields.
    • Applied new Apiezon N thermal grease to bottom flange of inner shield.
    • Re-clamped inner shield to cold plate using a dog clamp array.
    • Placed outer shield in contact with outer shield; rocking percieved when pressing on one quadrant of the annulus.
    • Confirmed continuity of shield RTDs.
  • Installed test mass, still with heater mounted to planar surface as in QIL/2715.
    • Confirmed suspension wires were contacting bare silicon OD surface in middle.
    • Confirmed continuity of heater and of workpiece RTD.
  • Closed up, pumped down, cooled down.
    • Roughing pump on at ~4:20 pm
    • Turbo on with no issues after 15 minute programmed delay
    • Cooling started at 5:15 pm, pressure was a few E-4 torr; no need to wait so long, that was just when we got around to it.

All images are (or soon to be) posted to the QIL Photo Dump.

Attachment 1: D2000310_y-004_section_double_linkage_20220207.png
D2000310_y-004_section_double_linkage_20220207.png
  2716   Fri Feb 4 14:00:19 2022 RadhikaDailyProgress 31 Jan Fastest Radiative Cooling run started

Attached are best fits for cooldown runs on 01/14 and 01/31. The setup for both cooldowns can be found in the previous ELOGs. We noticed that the outer shield did not cool as significantly on 01/31 than on 01/14, hinting that there might have been more thermal contact between the outer shield and cold plate / copper bar. 

The model considers the resistances of the following conductive elements (and uses these resistances as fit parameters):

   1.) Joint between cold head and copper bar
   2.) Bulk of copper bar
   3.) Joint between copper bar and flexible strap
   4.) Bulk of flexible strap
   5.) Joint between flexible strap and cold plate

These additions helped the model more closely resemble our recorded data, with a few exceptions:

   - At early cooldown times, the model seems to be underestimating the heat load on the inner shield and outer shield.

   - The best fit was performed on the inner shield and outer shield data (to fine tune elements of the cold linkage), so the test mass fit is not optimized. (This will be performed next to refine emissivity predictions.)

To identify bottlenecks in the cold linkage, I used the 01/31 model and tweaked the resistances to see which would provide the largest gains in cooldown. The results from such tweaks are below:

   - Halfing the resistance of the greased joints (1, 3, 5) made negligible change to the cooldown. 
   - Halfing the resistance of the bulk of the copper bar made significant improvement to cooldown (Attachment 3).
   - Halfing the resistance of the bulk of the flexible strap made some improvement to cooldown (Attachment 4).

From these observations, it seems like the greased joints are thermally efficient, and the bulk area of the copper bar appears to be the largest bottleneck. 

 

Attachment 1: Cooldown_0114_analyzed.pdf
Cooldown_0114_analyzed.pdf
Attachment 2: Cooldown_0131_analyzed.pdf
Cooldown_0131_analyzed.pdf
Attachment 3: Cooldown_0131_copperbar_halfR.pdf
Cooldown_0131_copperbar_halfR.pdf
Attachment 4: Cooldown_0131_flexstrap_halfR.pdf
Cooldown_0131_flexstrap_halfR.pdf
  2715   Mon Jan 31 15:40:21 2022 StephenDailyProgress 31 Jan Fastest Radiative Cooling run started

As discussed during the 21 Jan 2022 meeting, the next cryostat run will seek the fastest radiative cooling (again, see QIL/2706) through the following configuration choices:

  • move heater to test mass
  • add thermal grease to inner shield - coldplate interface (clamped joint)

Actions completed 27, 28, and 31 Jan 2022

  • Vent.
  • Remove Test Mass in frame.
    • For simplicity, we snipped the RTD lead (we didn't want to have to redo the cryovarnish joint).
      • Of course, the joint failed during the repairs so we had to redo the workpiece RTD cryovarnish joint anyway!
  • Use cryovarnish to affix heater.
    • Removed aquadag in area of joint using IPA-soaked Alpha wipe and scrubbing motion.
    • First, varnished cigarette paper anchor pad, then added heater.
    • We allowed the cryovarnish to cure overnight between the 27th and 28th, then allowed the joint to hang in its vertical orientation over the weekend to confirm its integrity.
  • Use conventional pin-socket arrangement to add junction to heater leads and workpiece RTD leads.
    • Confirm all RTD leads are functional, and repair crimp joints and Kapton tape insulation where needed.
  • Clean up aquadag flakes on coldplate.
  • Remove inner shield, add Apiezon M thermal grease, and bolt down again.
  • Touch up position of mylar shield on cold linkage, aluminum foil aperture covers.
    • All apertures are covered except:
      • Inner Shield: cold linkage (not an issue - partially occluded by copper bar, mylar tube extends into cold head view and mylar mitten)
      • Outer Shield: electrical port, cold linkage (aperture has been extended to have additional area underneath the original circular aperture)
  • Insert Test Mass in frame. Dog clamp down frame.
  • Pump down, cool down.
    • Pump down started at ~15:15.
    • Cool down started at 15:44.

Model updates required to reflect new configuration:

  • estimate conductive leak from 24 AWG heater leads to test mass
  • model conductance of cryovarnish joint between heater and test mass - does cigarette paper make a difference?
  • model greased joint of inner shield - cold plate interface
    • confirm clamp load is adequate for effective greased joint, at room temp and cold temp!
  2714   Fri Jan 28 10:31:15 2022 ranaSummaryTutorial videooh no, stap the units madness, aaaaahhhh!!! noooooo!!!!

What if I told you SI-units is way harder to miscalculate with - Matrix  Morpheus | Make a Meme

 

  2713   Tue Jan 25 14:01:27 2022 StephenDailyProgressCryo vacuum chamberHeat transfer between grease joints and pressure joints

There's also a convention to write "50 kgf" to designate "kilograms of force" (implying the same conversion Rana describes, multiplying by g). I see kgf enough in the mechanical engineering world that I wouldn't have been confounded, so I wanted to pass that along.

Quote:

I think its least confusing to just replace 50 kg g with 500 N. Writing 50 g can be misleading, it seems like 50 grams.

 

 

  2712   Tue Jan 25 13:15:40 2022 ranaDailyProgressCryo vacuum chamberHeat transfer between grease joints and pressure joints

I think its least confusing to just replace 50 kg g with 500 N. Writing 50 g can be misleading, it seems like 50 grams.

 

  2711   Tue Jan 25 10:24:14 2022 ranaDailyProgressCryo vacuum chamberHeat transfer between grease joints and pressure joints

For one surface resting on another, the force of the contact is the grativational force of the top surface. There's an implicit factor of g that cancels out from the ratio, so it becomes a ratio of masses. The heat conductance listed for a 50 kg Al-Al pressure joint is interpreted as the heat conductance for a force of 50g.

Quote:

Is there a units issue? 50 kg is a mass, but not a force.

 

  2710   Fri Jan 21 17:32:40 2022 ranaDailyProgressCryo vacuum chamberHeat transfer between grease joints and pressure joints

Is there a units issue? 50 kg is a mass, but not a force.

  2709   Fri Jan 21 13:28:48 2022 RadhikaDailyProgressCryo vacuum chamberHeat transfer between grease joints and pressure joints

I took a deep dive into Ekin 2.6 to understand heat tranfer between various joint types, specifically pressure joints and grease joints. This was motivated by the fact that modeling of the cold plate and inner shield temperatures seemed to be missing some key physics. 

Heat conductance through grease joints is dependent on the contact area of the surfaces:

\frac{\dot{q}}{\Delta T} = \frac{\dot{q}}{\Delta T}(1 cm^2) \frac{A}{1 cm^2}, where \frac{\dot{q}}{\Delta T} = the heat conductance. The heat conductance of a grease contact with area 1 cm2 can be found in Attachment 1, indicated by the region between the blue and red boxes (our temperature range).

Heat conductance through pressure joints is dependent on the force holding the surfaces in contact:

\frac{\dot{q}}{\Delta T} = \frac{\dot{q}}{\Delta T}(50kgf) \frac{F}{50kgf}, where the heat conductance of a pressure contact with force of 50kgf is found in Attachment 1, indicated by the orange box. (Megastat pressure contacts are between the cold plate and shields, so the Al-Al contact was referenced.)

[Added] Heat conductance through varnish joints are similar to grease joints, in that the conductance scales with contact area of the joint. (It follows the same formula as above for grease joints.) The heat conductance of a varnished contact with area 1 cm2 can be found in Attachment 1, and note that it is higher than that of greased contacts.

My efforts this week went into incorporating these heat links into the model. This required splitting up components into various parts (verging on a finite-element approach), since every joint is considered in addition to the elements themselves. A few notes:

1. For now, I assume the force between the inner shield and cold plate is 50 kg, for simplicity. Therefore, the heat conductance being used for this pressure joint is 4.5 W/K (from chart), and I am using this constant value across temperatures until a better estimation can be found.

2. For the grease joints, I estimate the heat conductance at 1 cm2 to be ~ 1 W/K, for the 50-300K range.

3. The contact between the outer shield and cold plate is not actually uniformly touching, as noted in 2706. I am not sure how to estimate the actual force between the surfaces, so I will add this as a fit parameter in the model. 

The new model with this incorporation still needs some adequate debugging, but I felt these were vital steps to ensure we get realistic answers regarding cooling power of the copper bar vs. LN2. Once I feel the model can be trusted, I will follow up with analysis of the new optimized cooldown. 

Attachment 1: thermal_conductance_joints.png
thermal_conductance_joints.png
  2708   Tue Jan 18 16:11:35 2022 StephenDailyProgressCryo vacuum chamberStatus of Fastest Radiative Cooling run started 14 Jan

I ended the first cooldown of the run at 12:49 today, approximately 92 hours after it had started. Workpiece temperature was 100 K - more data and model fitting will reveal more insights on the results.

I started a 5 W injection through the heater (dog-clamped to baseplate with indium gasket), which I intend to leave running until steady state temperature is reached, at least overnight. This power level will not present a risk to the system, so I feel comfortable with this static power input even though it is not interlocked.

Update after 25 hours [ATTACHMENT 1]

The 5 W heating of the shields has not quite leveled off yet, and the workpiece temperature resumed falling, showing the temperature had not yet reached a steady state. I will leave 5 W on until I see a steady state, then I will plan to turn off 5 W and allow the workpiece trend to level off fully. Summary:

  • with no heater input, after 90+ hours the workpiece was slowly cooling at a rate of ~ 1 K per 6 hours, and the inner shield temperature was stable at about 76 K.
  • with 5W heater input, after about 25 hours the workpiece is (still) slowly cooling at a rate of ~ 1 K per 18 hours, and the inner shield temperature is about 88 K and slowly rising about 0.1 K per 7 hours

See attachment 1 for data viewer screenshot which reflects the above summary.

Update after 145 hours [ATTACHMENT 2]

Continued injecting 5 W for the whole weekend. Over the last 24 hours, the workpiece temperature seems to have leveled off into a new cooling rate, while both shields alternated between heating and cooling. I think this can be described as a steady-enough state that I'm ending the cooldown.

Now, I'll let the chamber come up to temperature with the help of the 5 W load.

Attachment 1: 5w_input_qil_cryostat_20220119.png
5w_input_qil_cryostat_20220119.png
Attachment 2: 5w_input_121_to_145_hours_qil_cryostat_20220124.png
5w_input_121_to_145_hours_qil_cryostat_20220124.png
  2707   Mon Jan 10 16:46:19 2022 PacoMiscEquipment transfer2 inch GYRO optics

[Paco, Yehonathan]

This afternoon, we entered the QIL to look for a 2 in optic from the old GYRO experiment that might work as the PR2 in the BHD upgrade in 40m. We found some in the right most cabinet near where MEGASTAT is. We ended up taking two pieces (see Attachment #1 for evidence) one labeled "COATED 2"dia F.S. SUB S1: HR R>99.99% @ 1064NM S2: AR R<0.15% S/N:" and the other one from ATFilms, labeled "2"dia x 0.375" thk FS HR/AR @ 1064nm Run#s': V6-704 & V6-705" and a datasheet which specifies T = 0.02 % or 200 ppm.

Attachment 1: PXL_20220111_004553530.jpg
PXL_20220111_004553530.jpg
  2706   Fri Jan 7 14:54:58 2022 StephenDailyProgressCryo vacuum chamberFastest Radiative Cooling run started 14 Jan

As discussed during the 07 Jan 2022 meeting, the next cryostat run will seek the fastest radiative cooling through the following configuration choices:

  • eliminate radiation leak apertures
  • maximize emissivity of test mass volume
  • improve conductivity to outer shield

Actions completed 14 Jan 2022

  • Vent
  • Cover up aperture to viewport with foil (on both shields).
  • Remove foil from inner shield inner surface.
  • Place outer shield directly onto cold plate (no clamping force).
    • A bit unsatisfying, because the outer shield is not flat relative to the cold plate, but the shield is resting quite firmly with only a slight wobble when pressed. As with any plane-plane contact, there are a limited number of true points of contact, and the contact area is increased somewhat by foil that is folded around the outer shield or the cold plate.
      • I would estimate the contact area at ~5% of the total area of the bottom flange.
    • 2x G10 spacer washers were dropped into the lower volume whild making this switch.
    • Note that bottom lid of outer shield is still not conductively cooled, and therefore will be approximately room temperature.
  • Tidy up aluminum foil collar around bottom lid of outer shield.
  • Tidy up mylar shield around conductive link after inner shield
  • Pump down, cool down.
    • Pump down started at ~15:35.
    • Cool down started at 16:10.
  2704   Tue Dec 21 15:33:39 2021 StephenDailyProgressCryo vacuum chamberRadiative Cooling of Si Mass, with worse inner shield inner surface emissivity - next run repeating

I opened the QIL cryostat today for a health check and visual inspection before the next run. Because I saw a couple of interesting issues, I decided to redo the same run with more attention to detail on the closeout. I'm worried the outer shield may have been linked to the inner shield and baseplate enough to affect comparability with the prior run.

issues with this  run, requiring redo:

  • RTD wire for outer shield was clamped under lid of inner shield - this might have created a conductive link between the inner shield and the outer shield
  • outer shield was more wobbly than usual, and could have possibly been off of the three spacers - this might have created a conductive link between the outer shield and the baseplate

run details:

  • pumping started at 3:45 pm
  • turbo started with 15 minute delay at 20 torr
  • cryocooler started at 4:15 pm with active ion gauge pressure at 3 e-4 torr.

And since I didn't get to implement any of the intended next runs, here are some notes on other future runs of interest:

  • Si mass covered with Al foil (matte side facing out) - interested in making the emissivity of the test mass equal to that of the inner shield in the new config, for modeling.
    • (of course, this emissivity equivalence would be an approximation, as there is a large area of the test mass which is bare silicon)
  • outer shield clamped/resting on baseplate - this is predicted by Koji to be the most efficient cooling configuration.
  • shielding attached to structure holding Si mass (electropolished aluminum, aquadag aluminum, bare aluminum surfaces are all available.
  2703   Thu Dec 16 17:57:15 2021 Radhika, StephenSummaryCryo vacuum chamberMegastat geometric parameters

This ELOG serves as a compilation of known/measured geometric parameters of Megastat. This is informative for thermal modeling of the system, so I wanted to create a centralized reference. A reference to these dimensions has been added to the Wiki page

Chamber specs
   Outer Radius = 0.3048 m (12")
   Wall thickness = .00477 m (.188")
   Height = 0.3048 m (12")
   Flange thickness = .0254 m (1")

Outer shield specs
   Outer Radius = 0.2794 m (11")--> CAD .265 m (10.433")
   Height = 0.2286 m (9") --> CAD .206 m (8.110")
   Thickness = 2.90 mm (0.114") (CAD nominally 3 mm, but 9 gauge aluminum is standard)

Inner shield specs:
   Outer Radius = 0.244983 m (9.645") --> CAD .225 m (8.858")
   Height = 0.205994 m (8.11") --> CAD .192 m (7.559")
   Thickness = 2.90 mm (0.114")

Cold plate specs:
   Radius = CAD .254 m (10")
   Thickness = CAD .01498 m (.5897")

Test mass specs: (confirmed)
   Radius = 0.0508 m (2")
   Length = 0.1016 m (4")

Copper bar specs:
   Length = 0.508 m (20") --> note that center to center length is .440 m in CAD
   Width1 = 0.03175 m (1.25") (bulk cross section, could be approximated accross full length)
   Width2 = 0.049784 m (1.96") (cross section at cold head bolting interface)
   Thickness = 0.011684 m (0.46")

Coldhead specs:
   Radius = 0.03175 m (1.25")
   Thickness = .0516 m

CAD (.EASM) is located at https://caltech.app.box.com/folder/131056505764 (File path: Voyager > Mariner > CryoEngineering). Screenshot of current state is added as Attachment 1.

CAD (source file, .sldasm - SolidWorks 2021) may be accessed via the PDM Vault (File path: llpdmpro > voyager > rnd qil cryostat)

Attachment 1: D2000310_y-003_20211222.png
D2000310_y-003_20211222.png
  2702   Thu Dec 16 15:54:44 2021 Stephen, RadhikaDailyProgressCryo vacuum chamberRadiative Cooling of Si Mass, with worse inner shield inner surface emissivity - CTC100 temperature control success

This post will host plots and trends from this radiative cooling run. At a glance, the tuned CTC100 PI control was able to control the workpiece steady state temperature in this radiative cooling test within .005 K.

Run description: At 4 pm Wednesday, the workpiece temperature was at steady state from the QIL/2701 cooldown, a little less than 120 K. From 4pm Wednesday thru 5pm Thursday (25 hours) the CTC100 controller was actuating on the workpiece RTD temperature (cryovarnished to the suspended Si mass) using the resistive heater (dog clamped to baseplate with indium foil gasket). The conductive heating of the cold plate, and therefore the inner shield, led to radiative heating capacity (via ΔT)  that actuated on the temperature of the suspended test mass. As found in QIL/2643, the suspended Si mass is well isolated from conduction to the cold plate.

Before the run, the CTC100 PID controller was allowed to autotune using a long lag (600 s) and a moderate acutation step (10 W). After autotuning, the D term was still 0, which seemed fine.

Data: Attachment 1 plots cooldown curves for all RTDs during this run. Attachment 2 compares this run's test mass and inner shield temperature curves to those from the previous run (Aquadag on inner surface of inner shield). The expected result of this change (coating inner surface of inner shield with Al foil) is a weakened radiative coupling between the inner shield and test mass, leading to less effective cooling of the test mass. 

Initial observations from data:

1) The cold head temperature curve again suggests 2 time constants, and cooldown is identical.

2) The inner shield's cooldown is roughly unchanged.

3) The outer shield's temperature drops significantly more, indicating a stronger coupling to the inner shield. We will check for a conductive short the next time we open up.

4) The test mass's cooldown matches expectations (weaker radiative coupling).

[WIP - The data will be fitted and discussed]. More detailed analysis from fit to come, including from heater runs.

 

Attachment 1: cooldown_12-10_all.pdf
cooldown_12-10_all.pdf
Attachment 2: cooldown_12-10_vs_11-16.pdf
cooldown_12-10_vs_11-16.pdf
  2701   Fri Dec 10 15:58:57 2021 StephenDailyProgressCryo vacuum chamberRadiative Cooling of Si Mass, with worse inner shield inner surface emissivity

Started a new run this afternoon, with the following goals:

   1) confirm that the first run (QIL/2695) went smoothly, by performing a visual inspection in the chamber while setting up for the first run.

      - kapton tape affixing inner shield RTD lead junctions to inner shield had fallen. These junctions were simply hanging - not ideal, but apparently not too harmful. Not likely to impact temperatures, in my opinion, but could have led to shorts or glitches in data.

      - all RTDs appeared to be fixed and well-contacted to surfaces

      - Everything seemed to be in good shape with the copper bar, no apparent issues

   2) obtain a second run with similar configuration, now that the rigid copper bar linkage has been implemented.

   3) vary a single important parameter relative to the first run, namely the inner surface emissivity of the inner shield, so that the impact of that parameter may be observed.

      - Added Aluminum foil (matte side visible) to the inner shield inner surfaces (lid and cylinder, both). Anywhere there was previously black Aquadag, there is now matte aluminum foil.

      - Kept the same apertures for viewport access and for electrical and thermal connection passthrough, basically attempted to achieve identical shield coverage.

      - There is one small sliver of black aquadag visible at the location of the electrical leads, but I didn't worry about patching that small area.

Run Details:

   - Pumps on at  ~3:40 pm

   - Cooling started at 4:13 pm (pressure ~6 mTorr, rapidly falling with turbo pump spinning up from ~70% to ~85% over a 1 minute interval). Coldhead RTD is responsive.

   - All photos will be posted to the QIL Cryo Vacuum Chamber photo album.

   - Note from check in on Monday afternoon, ~ 69 hours after start: everything looks good, and the workpiece temperature (~127 K) seems to reflect the emissivity change.

  2699   Sun Dec 5 17:33:50 2021 ranaDailyProgressCryo vacuum chamberTransitioning MS cooldown fitting to MCMC scheme

thats looking good

You should try to use either the corner or getdist packages to plot the 'corner' plots commonly used when showing correlated posteriors (cf. https://emcee.readthedocs.io/en/v2.2.1/user/line/) so that we can see what's up with the other uncertainties

Quote:

This week I have been working towards a Markov-chain Monte Carlo (MCMC) approach for fitting Megastat cooldown data and obtaining estimates on various emissivity values. I started with a simple model, only simulating the radiative cooling from the inner shield to the test mass. I supply the test mass and inner shield temperature data, and the emissivity of Aquadag (coating both the TM and inner surface of IS) as the only fit parameter. I am using Stan for the modeling, and Attachment 1 is a copy of my stan model.

The results of the simple model are in Attachment 2. The emissivity of Aquadag is estimated as a gaussian centered around 0.7. I am still determining whether this result is "real", or if the simulation is simply returning my prior. I will look into this more before adding complexity to the model.

 

  2698   Fri Dec 3 09:31:40 2021 RadhikaDailyProgressCryo vacuum chamberTransitioning MS cooldown fitting to MCMC scheme

This week I have been working towards a Markov-chain Monte Carlo (MCMC) approach for fitting Megastat cooldown data and obtaining estimates on various emissivity values. I started with a simple model, only simulating the radiative cooling from the inner shield to the test mass. I supply the test mass and inner shield temperature data, and the emissivity of Aquadag (coating both the TM and inner surface of IS) as the only fit parameter. I am using Stan for the modeling, and Attachment 1 is a copy of my stan model.

The results of the simple model are in Attachment 2. The emissivity of Aquadag is estimated as a gaussian centered around 0.7. I am still determining whether this result is "real", or if the simulation is simply returning my prior. I will look into this more before adding complexity to the model.

Attachment 1: MS.stan
functions {

  real coupling(real e_inner, real e_outer, real A_inner, real A_outer) {
    
    real E;
    E = e_inner*e_outer / (e_outer + (A_inner/A_outer)*(e_inner - e_inner*e_outer));
    return E;
  }
  
  real Cp_Si(real T) {
... 66 more lines ...
Attachment 2: stan_blackcoat_fit.pdf
stan_blackcoat_fit.pdf
  2697   Fri Nov 19 14:01:40 2021 Stephen, RadhikaDailyProgressCryo vacuum chamberRadiative Cooling of Si Mass, with better shield emissivity

[WIP]

11/16 was the first Megastat cooldown after exchanging the copper braid linkage for a copper bar. Attachment 1 compares the cooldown trends for the test mass, inner and outer shields, and cold head. The solid curves are the new cooldown trends (copper bar), and the faded dashed curves are the previous cooldown trends (copper braid).

Immediate observations:

    - The coldhead has a reduced heat load, and interestingly a second time constant governs cooldown from ~2-35 hours.

    - The inner shield time constant is reduced significantly, but the inner shield experiences a slightly greater heat load at steady state.

    - The test mass cooling is improved as expected, given inner shield cooldown.

    - The coupling between the outer shield and inner shield has increased, resulting in greater cooling of the outer shield. This could explain the added heat load to the inner shield. 

Attachment 1: comp_cooldown_728_cooldown_1116.pdf
comp_cooldown_728_cooldown_1116.pdf
  2696   Fri Nov 19 09:36:52 2021 AnchalMiscEquipment transferBorrowing two 2" optics

I came and looked around for a Y1S HR coated 2" optic. I found two ATFilms labeled optics with information of only their substrate. The coating info is encoded in the run# but I could not find a place for what it means. So I'm gonna take these and measure the transmission in 40m.

  • 2"x.375" thk FS, HR/AR @ 1064, Run #'s: V6-704 & V6-705
  • 2" x 0.375" PL/PL Run #'s: V2-2239 & V2-2242
Attachment 1: PXL_20211119_174146457.jpg
PXL_20211119_174146457.jpg
  2695   Fri Nov 12 14:31:38 2021 Stephen, RadhikaDailyProgressCryo vacuum chamberRadiative Cooling of Si Mass, with better shield emissivity

In this phase, we are working toward improving our setup with a rigid copper bar, and obtaining a new data point for our radiative cooling thermal models for a suspended silicon mass. Since the past cooling runs of a silicon test mass did not yet incorporate aquadag-painted shields, we wanted to obtain a new data point in the model (in other words, we painted the shields in QIL/2645, but the next test was a PD measurement, so this is the first silicon test mass measurment after shields were painted). The improvement to the thermal linkage, now using a rigid copper bar with higher conductivity (ref. QIL/2666), is a second variable being changed simultaneously in the spirit of improving the cooldown time.

Refer to the prior post (QIL/2694) for the bulk of the blow-by-blow of configuring the chamber to use the rigid copper bar linkage. This post will describe the mounting of the Si mass, and the pump down and cool down.

  • The silicon mass with Aquadag barrel was dropped into the existing frame, with the previous wire arrangement and with no particular requirement on position or orientation (just best effort centering and leveling). Adjustments were done chamberside as access was easier.
  • The frame was lifted into the chamber, with the hanging mass supported by auxiliary fingers, and placed in an available area. Since conductive cooling was not a dominant mode of heat transfer in this setup (ref. QIL/2647), clamping to the baseplate was simply a single dog clamp on each foot of the frame.
  • The cigarette paper was cryovarnished to the surface in the bare central position. Once the cryovarnish was set, the RTD was cryovarnished to the cigarette paper pad. No  strain relief or thermal anchoring considerations were implemented. RTD continuity was verified.
  • Lids were bolted down and shields were finalized (avoiding shorting to copper bar, making sure foil drapes covering apertures were well positioned, etc.
  • Vacuum pumps on at ~3 pm, cryocooler on at 3:30 pm. At 4 pm, things are still looking good!

Closeout photos will be posted to the QIL Cryo Vacuum Chamber photo album.

Attachment 1: IMG_2738.jpeg
IMG_2738.jpeg
Attachment 2: IMG_2741.jpeg
IMG_2741.jpeg
Attachment 3: IMG_2742.jpeg
IMG_2742.jpeg
Attachment 4: IMG_2745.jpeg
IMG_2745.jpeg
Attachment 5: IMG_2747.jpeg
IMG_2747.jpeg
  2694   Fri Nov 12 14:21:32 2021 Stephen, RadhikaDailyProgressCryo vacuum chamberUpgrade to Rigid Copper Bar, and assorted transitions from PD testing

This post describes upgrade efforts from 10 - 16 November, with the following goals:

 - introducing a solid copper bar thermal linkage

 - shifting the setup away from PD testing

 - preparing for the next test (radiative cooling of Si)

Here are some highlights of the effort:

  • While removing the shields we found contaminants had plated near the line of sight surfaces at the optical window and the electrical feedthrough [Attachment 1]. The film was removed by IPA wipe [Attachment 2] and was not evident in any other location (presumably these were the cold surfaces in line of sight, so they received the most contaminant, but there may be a thinner deposition throughout!).
  • In order to install the copper rod, we needed to cut out slots in the outer shield and inner shield. We used a reciprocating saw and held the piece stiffly on a table [Attachment 3] [Attachment 4].
    • We tried to use large snips, but that failed to provide enough cutting force, especially where it was necessary to use the tip to access into the flanged corner. We also damaged a few of the electrical leads to the feedthrough (formerly used to wire the OSEMs) - this will require attention at next opportunity.
  • The copper rod itself was found to have a few issues:
    • Outer surfaces were somewhat tarnished and greasy, so an 80 grit aluminum oxide paper was used to clean up all surfaces [Attachment 5]. The surfaces were then wiped with IPA using Alpha wipes.
    • Slot width was matching the long thermal strap instead of the short thermal strap, so a drill press was used to add cut away the correct areas of one pair of holes [Attachment 6]. Later, this modification required larger washers in the stack under the bolt head.
    • The cold head had a larger OD than designed, so we were unable to use the corner holes to mount the RTD. There was not enough space at the corners to host the nut or washer. Instead one of the bolt pattern holes was used to host the RTD stack [Attachment 7]
  • The installation of the copper rod required the following steps:
    • We documented the previous state of the table [Attachment 8].
    • We removed the inner shield, the outer shield, and the old thermal linkage.
      • To access the cold head side of the linkage, we had to remove the lower conflat of the T.
    • We installed a 6" conflat flange, double faced with 4" bore (Lesker p/n DFF600X400), between the cryocooler and the T. This spacer raised the bottom face of the cold head to the correct height, so that the rigid copper bar runs approximately through the center of the shield apertures.
      • This required an order of new bolts with 3" length, to squeeze the three flanges (cryocooler, spacer, T) that are now stacked together.
    • We bolted the rigid copper bar to the coldhead, yawing the cryocooler to match the conflat bolt hole orientation while also pointing the copper bar down the axis of the arm.
      • This used new vented screws (UC Components p/n C-412-A) which are also silver plated, except in the location borrowed for the RTD stack as mentioned above [Attachment 7]. The RTD stack included a nut to adjust spring compression independently from screw threading.
      • Apiezon N thermal grease was applied on both surfaces to improve thermal conductivity across this joint.
      • We initially forgot to reinstall the mylar sheet radiation shielding that had been removed from the area around the cold head and around the linkage. This required that we reopen the bottom conflat to install the coldhead mitten, and that we pitch the aluminum shields away from the rigid bar to allow the mylar sheet to be inserted from the inside.
      • We found that the coldhead RTD had failed during intial mounting efforts, and a new RTD (actually one that had been desoldered from the setup previously, removed from the workpiece on 2021.08.07 but found to have no issue) was soldered and attached to the cold head, under the spring clamp.
    • Each shield was reinstalled, including:
      • The newly cut slots were used to pass shields over the rigid copper bar.
      • Electrical cabling was threaded through the usual apertures.
      • The outer shield was positioned on the G10 spacers.
      • Rough alignment was completed based on clearance from the rigid copper bar, and line of sight to optical window.
    • Final touchups were implemented:
      • Aluminum foil covered unused outer shield apertures.
      • A small aluminum foil panel was placed underneath the rigid copper bar, to cover the slot in the inner shield.
      • Final clamping of the inner shield and the heater (with indium gasket) were completed.
      • Thermal strap was used to link the cold baseplate to the rigid copper bar, with a bolted joint at the copper bar and a dog clamp joint at the baseplate. Apiezon N grease was applied to all contact faces.

The rest of the installation effort is captured in the next log post QIL/2695, to partition the items relevant to the radiative cooling of the silicon mass.

The photos here (and others) are posted to the QIL Cryo Vacuum Chamber photo album.

Attachment 1: IMG_0352.JPG
IMG_0352.JPG
Attachment 2: IMG_0353.JPG
IMG_0353.JPG
Attachment 3: IMG_0371.JPG
IMG_0371.JPG
Attachment 4: IMG_0370.JPG
IMG_0370.JPG
Attachment 5: IMG_0343.JPG
IMG_0343.JPG
Attachment 6: IMG_0351.JPG
IMG_0351.JPG
Attachment 7: IMG_0358.JPG
IMG_0358.JPG
Attachment 8: IMG_0346.JPG
IMG_0346.JPG
Attachment 9: IMG_2734.jpeg
IMG_2734.jpeg
Attachment 10: IMG_2740.jpeg
IMG_2740.jpeg
  2693   Fri Nov 12 10:54:55 2021 ranaSummary2um PhotodiodesResults from JPL PD: A1-Test3

IT would also be good if you could plot the RMS noise around 10 Hz and 100 Hz as a function of the bias and temperature, so we can see what the trends are. And how about post the data and scripts to the elog so we can munge data later?

  2692   Fri Nov 12 08:21:22 2021 AidanSummary2um PhotodiodesResults from JPL PD: A1-Test3

[Aidan]

Here is the analyzed data from Test 3 of the A1 JPL PD.

Note about the QE measurement: 

  • HOM beam QE was performed during the warm up phase. Collimating lens was fixed and beam pointing was optimized at 100mA before each measurement. Likely that not all power was on PD but that distribution was constant throughout measurement. Therefore, good proxy for shape of QE response.
  • Manual QE was performed with 25mA current, optimized collimating lens position (and thus the beam size on the PD). The data corresponds to 8.0mm between the lens mount and fiber mount. The beam pointing was optimized before each measurement at 25mA.
  • QE projection scales the "HOM beam QE" result to the manual QE measurements to project out expected QE performance vs temperature

Dark current is the output from the Keithley scan - the vertical scale is correct in Amps [ignore question mark]

Dark noise spectra are included for different bias levels and at different temperatures. Stll need to add ADC noise floor for these plots.

Notes from Test 3

~/JPL_PD/data/A1_test3/README

6-Oct-2021: done with cryocooler off and temperature increasing

PREAMP GAIN = 1E3

SR560 gain = 500

LD temp set point = 20.2kOhm

 

 

 

Attachment 1: A1_test3_nominal_QE.pdf
A1_test3_nominal_QE.pdf
Attachment 2: A1_test3_darkcurrent.pdf
A1_test3_darkcurrent.pdf
Attachment 3: A1_test3_noise_spectra_rev.pdf
A1_test3_noise_spectra_rev.pdf A1_test3_noise_spectra_rev.pdf A1_test3_noise_spectra_rev.pdf A1_test3_noise_spectra_rev.pdf A1_test3_noise_spectra_rev.pdf
  2691   Wed Oct 27 10:18:33 2021 Aidan, StephenSummaryCryo vacuum chamberInspection of Megastat post 408K event and pumping timeline

[Aidan]

I've attached a timeline of our inspection this afternoon along with today's pumping timeline,

Here is a brief summary of observations from previous pumping timelines. Today's pump down is consistent with previously observed timelines.


 

2:17PM – Assessing the impact of the 408K event in the MegaStat

Coldhead reached 363K (90C)

Innershield reached 403K (130C)

 

2:24PM – Aquadag E service temperature (149C)

https://www.laddresearch.com/index.php/lanotattachments/download/file/id/40/store/1/

Maximum service temperature in air* : 300°F (149°C) 

*Service temperature under vacuum conditions is significantly higher. Contact Acheson for specifics. 

2:25PM – Bringing Megastat back up to air for initial inspection

2:37PM – chamber is at air

 

2:41PM – removing bolts

 

 

3:13PM – initial inspection looks normal. No elevated amount of black particulates found on surfaces – consistent with or less than the amount seen last time we opened.

Stephen detected faint smell different from last time (“campfire”?)

  • One RTD connector did delaminate Aquadag from the inner shield

3:14PM - Stephen reattaching the RTD wiring that had delaminated. I wiped up visible particulates with isoproponal soaked wipe.

 

3:21PM – putting lid back on

3:30PM – lid on. Screws in finger tight

 

3:35PM – screws tight – ready to pump

3:40PM – pumping station on

 

Time (minutes) Pressure (Torr) Notes
0 7.5E2 Pirani gauge initially
1 7.5E2  
2 7.5E2  
4 7.5E2  
5 7.5E2  
7 7.5E2  
9 7.5E2  
10 7.5E2  
11 7.5E2  
12 4.3E2 Gauge starts reading decrease
13 2.2E2  
14 9.8E1  
15 5.5E1 Turbo ON
16 2.9E1  
17 1.6E1 Turbo at 44%
18 9.4E0  
19 4.9E0  
20 1.8E0 Turbo at 58%
21 3E-2 Turbo at 70%
22 1.1E-3 ION gauge readings from here. Turbo at 91%
23 7.5E-4 Turbo at 100%
24 6.3E-4  
25 5.9E-4  
26 5.5E-4

 (Cryo-cooler normally turned on around this time)

Not in this instance though

27 5.0E-4  
28 4.59E-4  
29 4.26E-4  
30 4.05E-4  
     

 

Attachment 1: IMG_5407.jpg
IMG_5407.jpg
Attachment 2: IMG_5408.jpg
IMG_5408.jpg
Attachment 3: IMG_5409.jpg
IMG_5409.jpg
Attachment 4: IMG_5410.jpg
IMG_5410.jpg
Attachment 5: IMG_5411.jpg
IMG_5411.jpg
Attachment 6: IMG_5412.jpg
IMG_5412.jpg
Attachment 7: IMG_5414.jpg
IMG_5414.jpg
Attachment 8: IMG_5415.jpg
IMG_5415.jpg
  2690   Tue Oct 26 08:43:38 2021 AidanUpdateCryo vacuum chamberCTC100 temperature alarm and heater shutoff

Instructions on how to enable the alarm and heater shut off for the CTC100.

Status: This reports the status of the alarm. If LATCH is enabled, this must be manually set to OFF once it has been enabled.

Mode: Set to "Level"

Latch: Optional to set to "YES" if desired.

Output: Set to "Heater"

Max: Set to desired maximum temperature.

The attached photos show:

  • the menu where the settings are ALARM entered
  • the main display just before the alarm is enabled (at 300.350K with a 1s delay)
  • the main display just after the alarm is enabled - note that the Heater Output has been set to 0W.
Attachment 1: Screenshot_from_2021-10-26_08-42-57.png
Screenshot_from_2021-10-26_08-42-57.png
Attachment 2: IMG_5371.jpg
IMG_5371.jpg
Attachment 3: IMG_5396.jpg
IMG_5396.jpg
Attachment 4: IMG_5401.jpg
IMG_5401.jpg
ELOG V3.1.3-